### v1

@@ 2 Results
Analyzing the effect of weight quantization, jitter and packet drops: The quantization of weight bits decides the number of shaped queues and PCP bits encoding required in the Ethernet-based implementation of $\pi^{2}$ networks (Fig. 2A,C). As shown in Fig. 7A, quantizing the ResNet-9 model on CIFAR-10 to 8-bit and 6-bit precision has a negligible effect on accuracy, maintaining baseline performance near 94.2%. However, a significant degradation in accuracy is observed for a 3-bit network without fine-tuning. Notably, applying quantization-aware fine-tuning helps recover much of this loss for lower-precision networks. An accuracy above 90% can be achieved using 3-bit quantized $\pi^{2}$ networks on fine-tuning. Inducing random packet drops in $\pi^{2}$-NN is similar to adding a dropout layer in conventional networks. Events are dropped independently in every layer with a predefined probability. We analyze the results of an 8-bit quantized ResNet-9 network trained on the CIFAR-10 dataset. Spiking time jitter in $\pi^{2}$-NN is similar to activation noise in conventional deep neural networks [30]. We investigate the results of an 8-bit and 4-bit quantized $\pi^{2}$-ResNet9 network trained on the CIFAR10 dataset. The results are reported in Fig. 7B,C and the fine-tuning details are elaborated in the Methods section 4.5,4.6.
@@ 3 Discussion
In this paper, we introduce a $\pi^{2}$ neuromorphic computing paradigm that leverages interconnect primitives: programmable delay, time-based sorting, and event dropping for computation and memory. Forward inference therefore becomes signal propagation and scheduling in time rather than the conventional multiply–accumulate (MAC) arithmetic. Viewed this way, $\pi^{2}$ offers a practical route to implement physical AI systems that are scalable while leveraging computational primitives inherent in the physical interconnect substrate (e.g., networking switches in packet-switching networks, waveguide delay lines in photonic networks[49]). The key contribution for this work is to show how to re-purpose existing Ethernet traffic shaping protocols - namely the CBS and ATS protocols to model the $\pi^{2}_{K}$ neuron dynamics and synaptic operations. By mapping neural operations directly onto these communication protocols, we transform commodity interconnect hardware into an in-network neuromorphic processor. To demonstrate this we used an OMNET++ network simulator, where due to the limitation of the simulator, we were only able to demonstrate a 3-layer $\pi^{2}$-NN network with 3-bit synaptic delays which were encoded using PCP tags within the Ethernet frames. However, our modeling studies show that the $\pi^{2}$ paradigm can be scaled, where we showed minimal to no drop in accuracy as the number of model parameters increase. The full scaling implication can only be realized when the $\pi^{2}$-NN is implemented on a switching hardware, using commercial off-the-shelf (COTS) or custom chipsets, which we describe next.
@@
Figure 8: A) Architecture showing how the $\pi^{2}$-NN could be implemented on a high-speed packet-switching hardware platform. Each port of the switch supports a bandwidth of $B$ Gbps. Two ports are reserved for interfacing with external inputs and outputs, while the remaining ports are used for intra-core communication. Incoming events (represented by their corresponding source (Tx) addresses) are parsed and classified based on their priority (PCP code) and destination address (Rx address) by the ingress processing and traffic classification unit. The ATS mechanism is reinterpreted to emulate a $\pi^{2}$-synapse by introducing controlled delays, and the CBS mechanism is modified to replicate the spiking behavior of a $\pi^{2}_{K}$ neuron. These processing blocks can be cascaded within the switch to emulate $m$ neurons and their corresponding synapses at every output port. The output events—such as one generated at time $T^{1}$ from an input at $T^{0}$—can be recursively routed back to the input ports for continued processing. Here, the superscript $T^{v}$ denotes the processing order of the event. B) Trends in the performance projections for large-scale $\pi^{2}$ networks
@@
Ethernet switches as $\pi^{2}$ neuromorphic processors : Modern switches now support the IEEE 802.1Q protocol suite, enable VLAN tagging [50], and incorporate advanced packet processing capabilities alongside traffic management features [51, 52, 10], including the CBS and ATS primitives summarized in Fig. 2. Thus, an ethernet switch/router [53, 11, 54] by itself can be used as a viable hardware substrates for the efficient and scalable realization of the $\pi^{2}$ computing paradigm. This not only obviates the need for new and time-consuming hardware development but also leverages future advances in networking hardware technology, as discussed in section 1. Although the current standards do not yet support the combination of multiple traffic shapers on the same queue, extensive analysis of such ideas has been done in recent works [55, 34], which we leverage here for the proposed $\pi^{2}$ architecture.
@@
Figure 8A shows the architecture of the $\pi^{2}$ neuromorphic core in which a population of neurons communicate over a high-performance Ethernet switching fabric. The switch is provisioned with $H_{in}$ ingress ports and $H_{out}$ egress ports. Among these, one input and one output port serve as interfaces to external systems, while the remaining ports facilitate intra-core communication. Each port is shared by $m$ neurons and operates with a bandwidth of $B$ Gbps, enabling high-throughput event exchange. The arriving frames, represented by their source addresses, are parsed and classified by the ingress processing and traffic classification unit. Within the egress pipeline, each output port is shared by a pool of $m$ neurons and their corresponding synapses, modeled using the proposed CBS and ATS traffic shaping protocols. These model the $\pi^{2}$ computing paradigm, for event processing and communication. This pipeline is replicated in parallel across every output port to accommodate $n=H_{out}m$ neurons. The details of the implementation are elaborated in the supplementary section 5.
Hardware and energy scaling of $\pi^{2}$ architectures: Interestingly, the evolution of Ethernet switch ASICs has progressed along a promising trajectory, with switching capacity doubling approximately every two years—driven by alternating advancements in the number of SerDes lanes and the signaling rate per lane [11]. By implementing the proposed neuromorphic core (Fig. 8A) using current state-of-the-art interconnect modules and Ethernet switches integrated with high-bandwidth memory (HBM), the total system power is estimated to be around 1.7 kilowatts for simulating approximately 71% of a brain-scale network (assuming the human brain has around 100 trillion synapses [20, 56]). This configuration delivers a peak throughput of 0.71 peta operations per second (petaOps/s) and achieves an energy efficiency of 2.37 picojoules per operation (pJ/Op). In contrast, earlier neuromorphic simulations demonstrate significantly higher power consumption even at smaller scales. For instance, the IBM Blue Gene/L supercomputer consumed approximately 665 kilowatts to simulate just 1% of the brain-scale network [57]. Similarly, projections for neuromorphic systems based on Intel Hala Point [58], TrueNorth and SpiNNaker estimate power requirements of around 2.6 kilowatts, 4 kilowatts (considering only four 1-kW backplanes) and more than 25 kilowatts, respectively, for 1% brain-scale workloads [20, 21]. A detailed analysis of the power and performance projections of the discussed implementation strategy using current state-of-the-art Ethernet switches is presented in the supplementary section.
The current trends in the evolution of Ethernet switches indicate that the switch radix count, SerDes speed (bandwidth), and power consumption of the switch ASIC have increased by 8x, 10x, and 8x, respectively, in the past 12 years [11]. Extrapolating and projecting an 80x growth in Ethernet switch bandwidth in the next 12 years as well, we follow a similar analysis and compute the power numbers for the system with 4096 (ports) x 1 TBps (bandwidth), consuming 4096 Watts (512x8 - 1pJ/bit). Further, consistent development in co-packaged optics (CPO) technology has led to an interconnect power density down to 1 pJ/bit [59]. Also, with the improvements in the stacking technology and the advent of lower technology nodes, a much better energy efficiency of less than 1 pJ/bit [60, 61] can be achieved in HBMs. Using these numbers, the system can support 4860 trillion synapses (48.6 times brain-scale). Following a similar analysis, the total power consumption of the interconnect system can be computed to be around 27.8 kW. The system has the potential to achieve 48.6 petaOps/s operating with an energy efficiency of 0.57 pJ/operation. Considering a brain-scale simulation with these technological advancements, with 100 billion neurons and 100 trillion synapses, the total power of the interconnect system can be computed to be less than 570 watts. The trends in technological advancements and the projected performance metrics of $\pi^{2}$ networks is presented in Fig. 8B and discussed further in the supplementary section.
@@
Spatiotemporal encoding in $\pi^{2}$-NN : In $\pi^{2}$-NN, the event encoding scheme is spatiotemporal, that is the information is carried by the subset of neurons that activate in each layer and, within that subset, by their temporal order of spiking. Because synaptic delays allow a single neuron’s event to be delayed by multiple offsets, a given unit can participate in multiple contributing groups that drive the next layer (supplementary section 2). Selectivity, therefore, arises from spatial sparsity (which units participate) and temporal ranking (when they arrive). This joint code is inherently combinatorial, as the number of distinct patterns grows with both the set of active neurons and their permutations in time [36][18]. Further, as presented in Eq. 1, every $\pi^{2}_{K}$ neuron is activated by the earliest arriving K events. The identity of the K earliest arriving spikes provides a natural back-pointer, enabling complete reconstruction of the causal route. This equivalence yields two interesting aspects: interpretability (explicit paths) and computational sparsity (only contenders on the best path are evaluated). From a training perspective, the back-pointers make credit assignment hardware-friendly as the synaptic weight updates are sparse (only K paths for every contributing neuron have to be updated), and the update rule is simple too [29]. Thus, $\pi^{2}$ networks are ideal for on-chip training strategies where memory, bandwidth, and energy are scarce. A single knob, K, sets how many earliest arrivals each neuron exploits, thus controlling computational sparsity and performance. Specifically, in K=1 networks, the first output to spike is the globally shortest path through the layered network; no analog accumulation is required. In this regime, the network follows shortest-delay paths, maximizing sparsity and minimizing latency. The predicted class is simply the output that first registers an event, making K=1 especially attractive for extreme sparse computing and latency-sensitive applications. Even though in this work, we have not fully exploited polychronicity to improve the capacity of the $\pi^{2}$-NN, future research will be able to use the proposed training infrastructure to incorporate such an encoding. Also, in this work we showed how $\pi^{2}$ neural networks can be efficiently trained using a standard GPU infrastructure. However, the training time could be significantly reduced once a dedicated $\pi^{2}$ hardware is available. This is because GPUs have been optimized to execute standard multiply-accumulate (MAC) and matrix operations, where as the fundamental operations in a $\pi^{2}$ network are sorting and delays which, as we showed in this paper, can be efficiently executed on a switching and interconnect hardware. Future work will therefore focus on realizing a dedicated $\pi^{2}$ computing platform using COTS hardware or dedicated ASIC chipsets.

### v3

@@ 2 Results
Analyzing the effect of weight quantization, jitter, and packet drops: The quantization of weight bits decides the number of shaped queues and PCP bits encoding required in the Ethernet-based implementation of $\pi^{2}$ networks (Fig. 2A,C). As shown in Fig. 7A, quantizing the ResNet-9 model on CIFAR-10 to 8-bit and 6-bit precision has a negligible effect on accuracy, maintaining baseline performance near 94.2%. However, a significant degradation in accuracy is observed for a 3-bit network without fine-tuning. Notably, applying quantization-aware fine-tuning helps recover much of this loss for lower-precision networks. An accuracy above 90% can be achieved using 3-bit quantized $\pi^{2}$ networks on fine-tuning. Inducing random packet drops in $\pi^{2}$-NN is similar to adding a dropout layer in conventional networks. Events are dropped independently in every layer with a predefined probability. We analyze the results of an 8-bit quantized ResNet-9 network trained on the CIFAR-10 dataset. Spiking time jitter in $\pi^{2}$-NN is similar to activation noise in conventional deep neural networks [stanojevic2024]. We investigate the results of an 8-bit and 4-bit quantized $\pi^{2}$-ResNet9 network trained on the CIFAR10 dataset. The results are reported in Fig. 7B,C, and the fine-tuning details are elaborated in the Methods section 4.5,4.6.
Analyzing the effects of spiking sparsity: Motivated by the energy benefits of spiking sparsity [stanojevic2024], we introduce a sparse-spiking variant of $\pi^{2}$, denoted $\pi^{2}_{s}$, which reduces the spike count per neuron to fewer than one spike on average. From Eq. 6, the first $K$ earliest presynaptic events determine the postsynaptic spike times. For small $K$, many $T_{i}^{+}$ events contribute negligibly to $T_{j}^{+}$ and $T_{j}^{-}$, indicating that only a small subset of early arrivals meaningfully influences the computation. Accordingly, in $\pi^{2}_{s}$ we remove the $T_{i}^{+}$ terms from the candidate sets in Eqs. 7,8, and retain only the earliest $P$ events of $T_{i}^{-}$ ($d\geq P\geq K$). The single-ended output in Eq. 5, previously encoded as a differential pair, is now represented by a single event occurring at $T_{j}^{-}$. Only these $P$ earliest events propagate to the next layer, yielding a spiking sparsity of $P/d$.
Mathematically, the real-valued $d$-dimensional input X is encoded into a single event spike time and fed as input to $\pi^{2}-NN$ as follows:
$T^{0-}=[A-X]_{+}$ (11)
The synaptic weights (W) are encoded as differential synaptic delays represented as $W^{+}=[B+W]_{+},W^{-}=[B-W]_{+}$ for all layers. Here, A and B are arbitrary constants, and $_{+}=\max(0,.)$ is a ReLU operation that ensures that the information represented in time is non-negative. The output of a $j^{th}$ neuron in a $\pi^{2}$-NN is one spike/event denoted by its respective time of occurrence $T_{j}^{-}$. This occurrence is computed as follows:
$\tilde{T}^{+}_{j}=\frac{M}{K}+\frac{1}{K}\sum_{k=1}^{K}t_{[k],j}^{+}\ \text{and}\ \ \tilde{T}^{-}_{j}=\frac{M}{K}+\frac{1}{K}\sum_{k=1}^{K}t_{[k],j}^{-}\quad\text{where}$ (12)
$t_{[k],j}^{+}\in\left\{T_{[i]}^{-}+W_{ij}^{-}\,\middle|\,i=1,\ldots,P\right\}$ (13)
$t_{[k],j}^{-}\in\left\{T_{[i]}^{-}+W_{ij}^{+}\,\middle|\,i=1,\ldots,P\right\}$ (14)
The terms $t_{[1]},t_{[2]},\ldots,t_{[K]}$ represent the first $K$ earliest arrival times among the inputs defined in the sets. The term $T_{[i]}^{-}$ represents the $i^{th}$ earliest arriving presynaptic events. To obtain a single-ended output, the differential spike times are combined as follows, and a nonlinear activation function (ReLu) $\sigma(.)$ is subsequently applied:
$)T_{j}=\sigma\ (\alpha(\tilde{T}^{-}_{j}-\tilde{T}^{+}_{j}$)) (15)
$T_{j}^{-}=V+[A-T_{j}]_{+}$ (16)
Here, $\alpha$ is a hyperparameter, and V is a constant added to maintain causality. The post-synaptic event for the neuron $j$ occurs at a time $T_{j}^{-}$ and is propagated to the next layer for further processing. This inherent selectivity provides a natural mechanism for enforcing high spiking sparsity, motivating the design of sparse-$\pi^{2}$ ($\pi^{2}_{s}$) variants that reduce event volume while maintaining functional performance.
When compared to the original $\pi^{2}$ formulation, the $\pi^{2}_{s}$ paradigm at least reduces the communication traffic by 50%. Although the computation remains differential, only a single spike (the $T_{j}^{-}$ event) is transmitted forward, cutting communication traffic by half while preserving the temporal dynamics of the original $\pi^{2}$ formulation. The spiking sparsity reduces from 2 spikes/neuron in $\pi^{2}$-NN to $P/d$ ($\leq 1$) spikes/neuron. From the training perspective on GPUs, we can enable training of larger networks, as we now have to sort a much smaller set consisting of only P data points. Initially, we experimented with $(P/d)$ of 1 across all the layers of a ResNet-18 network and achieved an accuracy of 95.3% and 77.2% respectively on the CIFAR-10 and CIFAR-100 datasets. Further, inspired by the work in [stanojevic2024] (they propose a spiking sparsity of 0.3 spikes per neuron), we experimented with a sparsity ratio of 1 in the first layer (as d was already small enough) and 0.3 in the remaining layers and achieved an accuracy of 95.2% for CIFAR10 and 77.1% for CIFAR100 datasets (0.7% drop from SOTA [r18]). The results of the CIFAR datasets trained on the $\pi^{2}_{s}$ based ResNet-18-based network (11.2M parameters) with this algorithm are presented in Fig.7d. The training details are elaborated in the Methods section 4.7.
@@ 3 Discussion
In this paper, we introduce a $\pi^{2}$ neuromorphic computing paradigm that leverages interconnect primitives: programmable delay, time-based sorting, and event dropping for computation and memory. Forward inference, therefore, becomes signal propagation and scheduling in time rather than the conventional multiply–accumulate (MAC) arithmetic. Viewed this way, $\pi^{2}$ offers a practical route to implement physical AI systems that are scalable while leveraging computational primitives inherent in the physical interconnect substrate (e.g., networking switches in packet-switching networks, waveguide delay lines in photonic networks[photo]). The key contribution of this work is to show how to re-purpose existing Ethernet traffic shaping protocols - namely the CBS and ATS protocols to model the $\pi^{2}_{K}$ neuron dynamics and synaptic operations. By mapping neural operations directly onto these communication protocols, we transform commodity interconnect hardware into an in-network neuromorphic processor. To demonstrate this, we used an OMNET++ network simulator, where, due to the limitation of the simulator, we were only able to demonstrate a 3-layer $\pi^{2}$-NN network with 3-bit synaptic delays, which were encoded using PCP tags within the Ethernet frames. However, our modeling studies show that the $\pi^{2}$ paradigm can be scaled, with minimal to no drop in accuracy as the number of model parameters increases. Notably, 3–4 bit effective precision is now consistent with frontier models, and the era of sub-2-bit LLMs underscores that meaningful computation does not require high numerical precision [llm]. However, higher precision bits can be incorporated using different strategies - a) We can increase the bits associated with the PCP field of the VLAN tag (4 bytes long). This will increase the number of shaped queues required for ATS, b) hierarchical delay decomposition, in which memory-access latency/routing delay provides coarse delay resolution while ATS-based queuing provides fine-grained scheduling, together enabling more than 8 bits of effective precision (as presented in 3); these approaches preserve throughput by retaining the pipeline structure of the $\pi^{2}$ neuron and synapse while allowing the effective synaptic precision to scale with minimal changes to the underlying switch microarchitecture. While validating these dynamics on hardware is ultimately desirable, access to commercial TSN platforms is limited and costly and lacks programmability, as documented in prior work [tsn_hw_2]. Notably, delay-based computation is a common feature in hierarchical Address Event Representation (AER) implementations [gert] and is supported on neuromorphic chips such as Loihi [liohi_delay], SpiNNaker [spin_delay], and TrueNorth [TN_delay]. Temporal sorting, causality, and buffering are inherent features of AER, asynchronous NOCs [temp, aer_sort]. Furthermore, in such architectures, constraining the buffer size makes event dropping an inherent feature. Credit accumulation is synonymous with membrane potential integration in the existing spiking neuron models, such as the integrate-and-fire (IAF) neuron [temp, iaf]. However, the full scaling implications can only be realized when the $\pi^{2}$-NN is implemented on switching hardware, using commercial off-the-shelf (COTS) or custom chipsets, as we describe next.
@@
Figure 8: A) Architecture showing how the $\pi^{2}$-NN could be implemented on a high-speed packet-switching hardware platform. Each port of the switch supports a bandwidth of $B$ Gbps. Two ports are reserved for interfacing with external inputs and outputs, while the remaining ports are used for intra-core communication. Incoming events (represented by their corresponding source (Tx) addresses) are parsed and classified based on their priority (PCP code) and destination address (Rx address) by the ingress processing and traffic classification unit. The ATS mechanism is reinterpreted to emulate a $\pi^{2}$-synapse by introducing controlled delays, and the CBS mechanism is modified to replicate the spiking behavior of a $\pi^{2}_{K}$ neuron. These processing blocks can be cascaded within the switch to emulate $m$ neurons and their corresponding synapses at every output port. The output events—such as one generated at time $T^{1}$ from an input at $T^{0}$—can be recursively routed back to the input ports for continued processing. Here, the superscript $T^{v}$ denotes the processing order of the event. B) Projected performance (petaOps/s) of $\pi^{2}$-based neuromorphic systems based on historical and extrapolated Ethernet switch trends, comparing two event-encoding granularities: (i) compact neuromorphic events consisting of $\log_{2}(n)$ bits, where n is the number of neurons per core, and (ii) standard Ethernet frames with a minimum size of 64 B.
@@
Ethernet switches as $\pi^{2}$ neuromorphic processors : Modern switches now support the IEEE 802.1Q protocol suite, enable VLAN tagging [ciscotsn], and incorporate advanced packet processing capabilities alongside traffic management features [broadcomm12.8, 25.6, marvel_tera], including the CBS and ATS primitives summarized in Fig. 2. Thus, an Ethernet switch/router [cpo, cisco, hpc] by itself can be used as an architecturally aligned substrate for the efficient and scalable realization of the $\pi^{2}$ computing paradigm. Although minor, protocol-aligned extensions/modifications to existing traffic-shaping mechanisms are required, $\pi^{2}$ can benefit from ongoing advances in interconnect and networking hardware technologies, as discussed in section 1. Although the current standards do not yet support the combination of multiple traffic shapers on the same queue, extensive analysis of such ideas has been done in recent works [atscbs1, atscbs2], which we leverage here for the proposed $\pi^{2}$ architecture. Further [tsn_comb] proposes a traffic shaping engine (TSE) for next-generation TSN-compliant NoC/SoC platforms, with generic hardware explicitly supporting the selection of different traffic-shaping mechanisms. Therefore, it can be anticipated that such capabilities will be available in the future.
@@
Figure 8A shows the architecture of the $\pi^{2}$ neuromorphic core in which a population of neurons communicates over a high-performance Ethernet switching fabric. The switch is provisioned with $H_{in}$ ingress ports and $H_{out}$ egress ports. Among these, one input and one output port serve as interfaces to external systems, while the remaining ports facilitate intra-core communication. Each port is shared by $m$ neurons and operates with a bandwidth of $B$ Gbps, enabling high-throughput event exchange. The arriving frames, represented by their source addresses, are parsed and classified by the ingress processing and traffic classification unit. Within the egress pipeline, each output port is shared by a pool of $m$ neurons and their corresponding synapses, modeled using the proposed CBS and ATS traffic shaping protocols. These model the $\pi^{2}$ computing paradigm for event processing and communication. This pipeline is replicated in parallel across every output port to accommodate $n=H_{out}m$ neurons. The details of the implementation are elaborated in the supplementary section 3.2 (Fig. S5).
Performance projections for large-scale $\pi^{2}$ architectures: The evolution of Ethernet switch ASICs has followed a sustained and favorable scaling trajectory, with switching capacity doubling approximately every two years due to alternating advances in SerDes lane count and per-lane signaling rate [cisco]. Over the past 12 years, switch radix, aggregate bandwidth, and total ASIC power have increased by approximately $8\times$, $10\times$, and $8\times$, respectively [cisco]. Extrapolating these trends suggests an additional $\sim 80\times$ increase in switch bandwidth over the next 12 years, enabling future-class switches with 4096 ports operating at 1 Tb/s per port at an estimated power of $\sim$4 kW (assuming $\sim$1 pJ/bit). In parallel, advances in co-packaged optics and 3D-stacked memory technologies have reduced interconnect and HBM energy efficiencies toward the sub-1 pJ/bit regime [cpoprogress, hbm_energy7, hbm_energy8]. Together, these trends point toward future systems dominated by bandwidth-rich, energy-efficient interconnect and memory fabrics.
Under the $\pi^{2}$ paradigm, these trends fundamentally alter the role of the interconnect: when computation is embedded directly within interconnect dynamics, the effective number of operations increases monotonically with available bandwidth (Fig. 8B), converting communication energy—traditionally treated as overhead—into useful computation. The associated technology trends and projected performance scaling of $\pi^{2}$ networks are discussed further in the supplementary section 3.2, 3.3 (Table S2, S3,S4). The performance projections presented here should be interpreted as architectural feasibility trends and idealized upper bounds for a dedicated $\pi^{2}$ deployment, in which a high-bandwidth switch datapath is reserved exclusively for $\pi^{2}$ event processing and integrated with external memory (HBM/DRAM). The analysis assumes sufficient routing capacity and memory bandwidth; in practice, memory capacity and access energy will impose additional constraints, particularly at brain scale. These limitations, however, are common to large-scale neuromorphic and HPC systems and are not unique to $\pi^{2}$. Accordingly, this work should be viewed as an architectural feasibility study rather than a claim about the capabilities of existing commercial Ethernet switches.
A plausible approach to overcoming the limited on-chip memory of individual Ethernet switches is to organize a $\pi^{2}$ system as a distributed compute–memory mesh of interconnected switches. This design is conceptually related to hierarchical Address-Event Representation (AER) architectures, which achieve scalability by distributing routing and neural state across multiple stages rather than relying on centralized memory [gert, aer_mem]. As in hierarchical AER, each switch maintains only local routing information and implements a limited subset of neurons and synapses. However, $\pi^{2}$ extends this model by explicitly repurposing routing latency as a computational primitive: events are forwarded using standard packet-switching mechanisms, while the number of hops traversed through the mesh encodes synaptic delay. In this way, network traversal directly implements neural timing rather than acting as communication overhead, enabling scalability by trading local memory capacity for network depth. This perspective aligns with the TEMP computational paradigm, which exploits intrinsic interconnect timing semantics for event-driven computation [temp], and suggests that interconnect hierarchy can serve as a viable substrate for large-scale neuromorphic computation.
What distinguishes $\pi^{2}$ from conventional systems is that, in today’s interconnect-centric platforms, Ethernet switches minimally contribute towards the effective computation, despite consuming a substantial fraction of total system power. By contrast, $\pi^{2}$ repurposes this otherwise passive communication energy to realize computation directly through interconnect dynamics. As a result, even when constrained by realistic packet sizes and memory bottlenecks, the interconnect transitions from an energy sink to an active computational substrate. This shift fundamentally alters the performance–energy trade-off, suggesting that future gains in interconnect bandwidth—driven by advances in SerDes, radix, and co-packaged optics—can be directly converted into increased effective computation under the $\pi^{2}$ model.
The architectural stance adopted in this work assumes a dedicated Ethernet switch datapath for $\pi^{2}$ processing consistent with other neuromorphic platforms, such as Loihi [liohi_delay] and Spinnaker[spin_delay], that also rely on dedicated interconnect and compute resources. In conventional neuromorphic substrates, scaling is limited by custom chip manufacturing and by challenges in memory–compute integration. However, this is not the case with Ethernet switches, whose capacity doubles every two years and is moving towards more energy-efficient processing. Thus, a dedicated switch deployment preserves the scalability advantages of networking hardware while enabling in-network computation. That said, multi-tenancy is feasible if required. Ethernet switches already support strong traffic-isolation mechanisms that separate best-effort and time-sensitive flows, including the use of distinct VLAN domains, reserved hardware queues [queue], and frame preemption strategies [fp] to guarantee QoS. These mechanisms could be leveraged to isolate $\pi^{2}$ traffic from conventional workloads, though performance would then depend on the level of resource reservation provided to $\pi^{2}$.
@@
Spatiotemporal encoding in $\pi^{2}$-NN : In $\pi^{2}$-NN, the event encoding scheme is spatiotemporal; that is, the information is carried by the subset of neurons that activate in each layer and, within that subset, by their temporal order of spiking. Because synaptic delays allow a single neuron’s event to be delayed by multiple offsets, a given unit can participate in multiple contributing groups that drive the next layer (supplementary section 2 (Fig. S2)). Selectivity, therefore, arises from spatial sparsity (which units participate) and temporal ranking (when they arrive). This joint code is inherently combinatorial, as the number of distinct patterns grows with both the set of active neurons and their permutations in time [poly][dendrocentric]. Further, as presented in Eq. 1, every $\pi^{2}_{K}$ neuron is activated by the earliest arriving K events. The identity of the K earliest arriving spikes provides a natural back-pointer, enabling complete reconstruction of the causal route. This equivalence yields two interesting aspects: interpretability (explicit paths) and computational sparsity (only contenders on the best path are evaluated). From a training perspective, the back-pointers make credit assignment hardware-friendly as the synaptic weight updates are sparse (only K paths for every contributing neuron have to be updated), and the update rule is simple too [mulp_mp]. Thus, $\pi^{2}$ networks are ideal for on-chip training strategies where memory, bandwidth, and energy are scarce.
A single knob, K, sets how many earliest arrivals each neuron exploits, thus controlling computational sparsity and performance. Specifically, in K=1 networks, the first output to spike is the globally shortest path through the layered network; no analog accumulation is required. In this regime, the network follows shortest-delay paths, maximizing sparsity and minimizing latency. The predicted class is simply the output that first registers an event, making K=1 especially attractive for extreme sparse computing applications. Also, the proposed formulation is closely related to established temporal coding strategies such as M-of-N, K-based, and rank-order spike encoding [mn, dendrocentric, furber]. As noted in prior work [temp], these encoding schemes are well-suited to neuromorphic hardware, offer higher information content per spike, and promote energy-efficient sparse computation by emphasizing temporal order rather than firing rates [dendrocentric, ttfs]. Even though in this work, we have not fully exploited polychronicity to improve the capacity of the $\pi^{2}$-NN, future research will be able to use the proposed training infrastructure to incorporate such an encoding. In this work, we showed how $\pi^{2}$ neural networks can be efficiently trained using standard GPU infrastructure. Training—whether distilled, or native—is a one-time cost, whereas inference is estimated to account for up to 90% of the total computational cost of large-scale AI deployments [aws]. Thus, even if teacher networks are used during the training process, the long-term computational and energy benefits of $\pi^{2}$ arise entirely during inference, where the architecture operates independently of any conventional hardware substrate. Further, the training time could be significantly reduced once dedicated $\pi^{2}$ hardware is available. This is because GPUs have been optimized to execute standard multiply-accumulate (MAC) and matrix operations, whereas the fundamental operations in a $\pi^{2}$ network are sorting and delays, which, as we showed in this paper, can be efficiently executed on a switching and interconnect hardware. This bottleneck reflects a constraint of current hardware, not a limitation of the $\pi^{2}$ learning rule. A dedicated $\pi^{2}$ accelerator would remove the GPU bottleneck entirely and enable large-scale training without relying on MAC-based networks. Thus, the observed training bottleneck is a technology limitation of current GPUs, not a fundamental restriction of the $\pi^{2}$ learning rule. Future work will therefore focus on realizing a dedicated $\pi^{2}$ computing platform using COTS hardware or dedicated ASIC chipsets.